At the beginning of the year, Young Company bought three used machines from Vince, Inc. The machines immediately were overhauled, were installed, and started operating. Because the machines were different, each was recorded separately in the accounts.

Machine A Machine B Machine C
Amount paid for asset $ 7,950 $ 26,300 $ 10,000
Installation costs 250 650 550
Renovation costs prior to use 2,200 1,550 2,050
Repairs after production began 470 420 500

By the end of the first year, each machine had been operating 8,000 hours.

Required:

  1. Compute the cost of each machine.
  2. Prepare the journal entry to record depreciation expense at the end of year 1, assuming the following: TIP: Remember that the formula for double-declining-balance uses cost minus accumulated depreciation (not residual value).
Estimates
Machine Life Residual Value Depreciation Method
A 5 years $ 200 Straight-line
B 40,000 hours 800 Units-of-production
C 6 years 2,000 Double-declining-balance
Answer #1

Requirement 1

Total Cost of Machine
Machine A $    10,400.00
Machine B $    28,500.00
Machine C $    12,600.00

Working

Amount paid for assets Installation cost Renovation cost Total Cost of Machine
Machine A $         7,950.00 $            250.00 $           2,200.00 $    10,400.00
Machine B $       26,300.00 $            650.00 $           1,550.00 $    28,500.00
Machine C $       10,000.00 $            550.00 $           2,050.00 $    12,600.00

Requirement 2

Transaction General Journal Debit Credit
1.. Depreciation expense $    11,780.00
           Accumulated depreciation-Machine A $        2,040.00
           Accumulated depreciation-Machine B $        5,540.00
           Accumulated depreciation-Machine C $        4,200.00
(To record depreciation expense)

Working

Straight line Method
A Cost $ 10,400
B Residual Value $ 200
C=A - B Depreciable base $ 10,200
D Life [in years left ]                                  5
E=C/D Annual SLM depreciation $               2,040.00

.

Units of Production method
A Cost $ 28,500
B Residual Value $ 800
C=A - B Depreciable base $ 27,700
D Usage in units(in Hours) 40000
E Depreciation per hour $                        0.69

.

Depreciation schedule-Units of Activity
Year Book Value Usage Depreciation expense Ending Book Value Accumulated Depreciation
1 $                28,500 8000 $                      5,540 $                    22,960 $                  5,540

.

Double declining Method
A Cost $ 12,600
B Residual Value $ 2,000
C=A - B Depreciable base $ 10,600
D Life [in years] 6
E=C/D Annual SLM depreciation $ 1,767
F=E/C SLM Rate 16.67%
G=F x 2 DDB Rate 33.33%

.

Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
1 $ 12,600 33.33% $ 4,200 $ 4,200 $ 8,400